Particular integral of (D^2 +3DD' + 2D'^2) = 12xy
1
Expert's answer
2021-07-19T05:42:13-0400
"(\ud835\udc37^ 2 +3DD'+ 2\ud835\udc37'^2)\ud835\udc67 = 12\ud835\udc65\ud835\udc66."

"\ud835\udc37^ 2 +3DD'+2 \ud835\udc37'^2=(D+D')(D+2D')"

"(\ud835\udc37^ 2 +3DD'+ 2\ud835\udc37'^2)\ud835\udc67 =(D+D')(D+2D')z"

"(D+D')(D+2D')z=12xy"

The auxiliary equation of the given equation is


"(k+1)(k+2)=0"

"k_1=-1, k_2=-2"

Than the complementary function of the given equation is


"u=\varphi_1(y-x)+\varphi_2(y-2x)"

Than partial integral


"P.I.=\dfrac{1}{(D+D')(D+2D')}(12xy)"

"=\dfrac{1}{D(1+\dfrac{D'}{D})D(1+\dfrac{2D'}{D})}(12xy)"

"=\dfrac{1}{D^2}\bigg[1+\dfrac{D'}{D}\bigg]^{-1}\bigg[1+\dfrac{2D'}{D}\bigg]^{-1}(12xy)"

"\dfrac{1}{1+t}=1-t+t^2-t^3+t^4-t^5+..., |t|<1"

"P.I.=\dfrac{1}{D^2}\bigg(1-\dfrac{D'}{D}+...\bigg)\bigg(1-\dfrac{2D'}{D}+...\bigg)(12xy)"

"=\dfrac{1}{D^2}\bigg(1-\dfrac{3D'}{D}+\dfrac{2D'^2}{D^2}+...\bigg)(12xy)"

"=\dfrac{1}{D^2}\bigg(12xy-\dfrac{3}{D}\dfrac{\partial}{\partial y}(12xy)"

"+\dfrac{2}{D^2}\dfrac{\partial^2}{\partial y^2}(12xy)+...\bigg)"

"=\dfrac{1}{D^2}\bigg(12xy-\dfrac{36}{D}x+\dfrac{2}{D^2}(0)+...\bigg)"

"=\dfrac{1}{D^2}\bigg(12xy-36\int xdx\bigg)"

"=\dfrac{1}{D}\bigg(\int(12xy-18x^2)dx\bigg)"

"=\dfrac{1}{D}(6x^2y-6x^3)=\int(6x^2y-6x^3)dx"

"=2x^3y-\dfrac{3}{2}x^4"

"P.I.=2x^3y-\dfrac{3}{2}x^4"

"z=\varphi_1(y-x)+\varphi_2(y-2x)+2x^3y-\dfrac{3}{2}x^4"

where "\varphi_1" and "\varphi_2" are arbitrary functions.